Sensor trouble

Submitted by admin on Fri, 07/20/2018 - 09:30

This page describes abnormal strong motion data caused by sensor troubles.

Station YKH

  • The N303°E components of the sensors on the second basement floor (B2F) and the 23rd floor (23F) at the station YKH were abnormal on records from March, 2010 to April, 2011. The 303-B2F and 303-23F components were filled with zeros. The trouble was recovered in July, 2011.

Station NMW

  • The N308°E component of the sensor on the east of basement floor (B1FE) at the station NMW has not been functioning since February, 2011. The 308-B1FE component (channel#8) of data files has strange signal.
  • The sensor on the east of basement floor (B1FE) has been repaired in June, 2011.

Station NGY

  • The up-down component of the sensor on the ground (GL) at the station NGY has not been functioning since February, 2011. The UP-GL component (channel#3) of data files has no signal.
  • The sensor on the ground (GL) has been repaired in June, 2011.

Station SND

  • The up-down component of the sensor 40 meters in the ground (G40) at the station SND has not been functioning since November, 2007. The UP-G40 component (channel#9) of data files has no signal.
  • The N074°E component of the sensor 40 meters in the ground (G40) at the station SND has not been functioning since October 2008. The 074-G40 component (channel#7) of data files has no signal.

Station ANX

  • The N180°E component of the sensor 14 meters in the ground A (A14) at the station ANX has not been functioning since July 2006. The 180-A14 component (channel#4) of data files has abnormal signal.
  • The N180°E component of the sensor 5F of the main building (M5C) at the station ANX has not been functioning since December 9, 2009. The 180-M5C component (channel#58) of data files has earthquake-like signal, but amplitudes are not reliable.
  • Signals of UP-A89 (channel#12) and 270-BFS (channel#29) are abnormal from November 1, 2013.
  • Signals of 180-A14 (channel#4), UP-A14 (channel#6), UP-A43 (channel#9) and 180-C01 (channel#19) are abnormal from November, 2013.

Station MYK

  • The sensors on the first floor (01F) and on the ground (GL) were damaged by tsunami caused by the 2011 Tohoku Earthquake. Some abnormal signals can be found records 30 minutes after the earthquake.
  • The the sensor on the ground (GL) is still out of order. Channels 7 to 9 have been removed from data files since May, 2011.
  • The sensor on the ground (GL) was reconnected in April, 2012. The strong motion files on and after April 12, 2012, have 9-channel data. But the sensor on the ground (GL) has trouble in signal insulation, so obvious noise can be observed on the records.
  • The N167°E component of The the sensor on the seventh floor (07F) was in trouble. The records of 167-07F (ch#04) from January 16 to February 2, 2013 are abnormal.

Station TKC

  • The sensor on the basement floor (B1F) was in bad condition from May 9 to June 8, 1012. Signals in low frequency range were unstable in all channels. The sensor was replaced on June 8, 1012 and the trouble was settled.

Station MNM

  • Signal of 270-01F(ch#5) is abnormal from November 29, 2013.

Station CG2

  • Signal of 208-B4F(ch#1) is abnormal from February 5, 2016.